OpenStudy (amtran_bus):

What is the first derivative of f(x)=2x^2+2x?

OpenStudy (anonymous):

f'(x) = 4x+2

OpenStudy (amtran_bus):

Can you walk me through the steps?

Parth (parthkohli):

\[\text D(2x^2 + 2x) = \text D(2x^2) + \text D(2x) = 4x + 2\]

OpenStudy (amtran_bus):

Thanks @ParthKohli Can you show me the steps with f(x)=c, which is 0?

Parth (parthkohli):

\[f(x) = c = cx^0\]Now the power rule.

OpenStudy (amtran_bus):

Yea, that is what I was thinking. Thanks!

Parth (parthkohli):

You're welcome!

OpenStudy (amtran_bus):

@ParthKohli , sorry to bother, but can you show me once more how the 2x is just 2 and not 2x? I use the forumla f(x+delta x)-f(x)/delta x

OpenStudy (anonymous):

2(1)x^1-1 2x^0 2(1) 2

OpenStudy (amtran_bus):

Thankss!